OpenStudy (anonymous):

How do I find the center and the radius of (x – 5)^2 + (y + 2)^2 = 16?

OpenStudy (anonymous):

center of \[(x-h)^2+(y-k)^2=r^2\] is \((h,k)\)

OpenStudy (anonymous):

and the radius is \(r\)

OpenStudy (anonymous):

in your example \(r^2=16\) so \(r\) is easy to find

OpenStudy (anonymous):

R would be 8, right?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

Wait I mean 4!

OpenStudy (anonymous):

\(8^2=64\) not \(16\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how about the center, you go that?

OpenStudy (anonymous):

Hm,not really..

OpenStudy (anonymous):

(h,k) = (5,2)?

OpenStudy (anonymous):

\[(x-\color{red}h)^2+(y-\color{blue}k)^2=r^2\] has center \((\color{red}h,\color{blue}k)\) close but not quite

OpenStudy (anonymous):

Oh so (h,k)=(2,5)? Sorry this is a new chapter for me..!

OpenStudy (anonymous):

\[(x-5)^2+(y+2)^2=4^2\] \[(x-5)^2+(y--(-2))^2=4^2\] center is \((5,-2)\)

OpenStudy (anonymous):

typo there, too many minus signs i meant \[(x-5)^2+(y-(-2))^2=4^2\] center is \((5,-2)\) and radius is \(4\)

OpenStudy (anonymous):

you were closer the first time. the \(h\) i.e. the first coordinate goes with the \(x\) and the \(k\) the second coordinate goes with the \(y\)

OpenStudy (anonymous):

Okay, got it. So the center is just an equation?

OpenStudy (anonymous):

Thank you so much, by the way! I really appreciate it! c: